OpenStudy (dehelloo):

What is the turning point of y= x^2+x-2

OpenStudy (bahrom7893):

Find the vertex

OpenStudy (anonymous):

if you want the vertex? y=x^2+x-2 y=(x^2+x)-2 y=(x^2+x+0.25-0.25)-2 move the -0.25 out of the brackets: y= (x^2+x+0.25)-2.25 factor: y=(x+0.5)^2 -2.25 vertex: (-0.5,-2.25)